# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ VII.2 Region 2: comparison with the local RN approximation
Figure VII.7: computed R vs. RN value of R
@@
In Figure VII.7a we compare the numerically-computed $R$ with that corresponding to the RN metric (in double null Eddington coordinates) but with its $M$ and $Q$ parameters replaced with $M(v)$ and $Q^{v}(v)$. Notice the complete overlap of the two curves .
Subsequently, Figure VII.7b shows $R(u)$ along the last $v$ ray, versus RN with constant $M,Q$ corresponding to $M(v=900)$ and $Q^{v}(v=900)$. Notice that because there is no drift of $M$ and $Q$ parameters in the $u$ direction, the two curves nicely overlap.
Figure VII.8: S computed vs. RN
@@
We move on to show $S$ in Figure VII.8 in region 2, versus the value of $S$ in the corresponding RN metric, determined by
$\frac{e^{S}}{r}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (VII.2)
In Panel VII.8a we plot $S(v)$ along the u=60u=-60 ray. Then in panel VII.8b we plot $S(u)$ along the last ray $v=900$. In both cases, the two curves are in good agreement.
VII.3 Region 3: comparison with the analytical approximation
In this subsection, we will compare the numerically computed quantities $R_{,u},$$R_{,v},$$S_{,u}$$S_{,v}$ in region 3 with their predicted values according to Eqs. (IV.5,IV.16,IV.40,IV.41).
Figure VII.9: Numerical result and analytical approximation for $S_{,v}$, $S_{,u}$
@@ VII.3 Region 3: comparison with the analytical approximation
We start with Figure VII.9a, which compares $S_{,v}$ and $\kappa_{-}(v)$, both displayed as a function of $v$. They are indeed very close, as predicted by the analytical approximation. Then we see in Figure VII.9b that the approximation is good for $S_{,u}$ as well, when comparing it to $\kappa_{-}^{u}$.
Figure VII.10: Analytical approximation of $R_{,v}$, $R_{,u}$
@@
Moving on to the derivatives of $R$, we see in Figure VII.10a that the analytical approximation Eq. (IV.40) works, comparing $R_{,v}$ with $\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ as they vary as functions of $v$.
Finally, we show in Figure VII.10b the quantities $R_{,u}$ and $\frac{\widetilde{T}_{uu}^{(-)}}{\kappa_{-}^{u}}$ in region 3. These quantities should both vanish, as mentioned in subsection II.8, because we have no sources in the present run. In the figure we see that both quantities are very small. It is remarkable that their erroneous non-vanishing value is still consistent with Eq. (IV.41). Our interpretation of this situation is as follows: these erroneous non-vanishing values happens because of finite-$du$ error. This causes a drift in $\widetilde{T}_{uu}$ leading to non-vanishing $\widetilde{T}_{uu}^{(-)}$– which in turn leads to non-vanishing $R_{,u}$ in region 3, in accordance with Eq. (IV.41).
To check this interpretation, we show in Figure VII.11 the magnitude of $R_{,u}$ for two different values of $du$: $\frac{1}{100}$ and $\frac{1}{50}$. The third curve shows the $du=\frac{1}{50}$ case multiplied by $\frac{1}{4}$. We can see that it perfectly overlaps with the curve of $du=\frac{1}{100}$. This shows that the erroneous non-vanishing value of $R_{,u}$ indeed scales with $du^{2}$, implying that it is associated with a finite-$du$ truncation error.
Figure VII.11: $R_{,u}$ with different $du$ values on last v ray
@@
VII.4 Deviation from the analytical approximation (region 3)
@@
First, we plot the deviation $S_{,v}+\kappa_{-}$ in Figure VII.12a. (To appreciate the magnitude of this deviation, note that it is roughly about 1% of the value of $S_{,v}$ as seen in Figure VII.9a.)
Figure VII.12: Deviation from analytical approximation of $S_{,v}$, $R_{,v}$
@@
We also show the deviation of $R_{,v}$ in Figure VII.12b
To analyze whether this deviation scales with $\varepsilon_{0},$we do two runs: one with $\varepsilon_{0}=\frac{1}{1000}$ as usual, and another one with $\varepsilon_{0}=\frac{1}{500}.$ To compare the two different runs, we plot the mentioned deviations as a function of $M^{v}(v)$ instead of as a function of $v$, to account for the fact that the evaporation rate is different in these two cases. Figure VII.13 compares two different $\varepsilon_{0}$ runs. (Note that in this plot, the low $v$ values correspond to high $M$, with $v=0$ at $M=1$.) If the deviation is proportional to $\varepsilon_{0}$, we would expect it to be twice as big in $\varepsilon_{0}=\frac{1}{500}$ compared to $\varepsilon_{0}=\frac{1}{1000}$.
Figure VII.13: $S_{,v}+\kappa_{-}$ with two $\varepsilon_{0}$ values, scaled by a factor of $\frac{1}{2}$, on the last u ray
To test this assumption, Figure VII.13 scales the curve of $\varepsilon_{0}=\frac{1}{500}$ by $\frac{1}{2}$ (green curve), and now we can see the two curves (green and blue) overlap, confirming that the deviation indeed scales with $\varepsilon_{0}$–meaning that it is a finite-$\varepsilon_{0}$ effect. We also display the unscaled $\varepsilon_{0}=\frac{1}{500}$ curve, as is, in orange.
We also wish to know how this finite-$\varepsilon_{0}$ deviation scales with $M$. From basic semiclassical scaling, one would expect this deviation to scale as $\propto M^{-3}$. This is because the dimensions of $S_{,v}$ is $S_{,v}]=[l]^{-1}$ (as $S$ is dimensionless), and we need to multiply this by the dimensionless factor $\frac{\hbar}{M^{2}}$ (as this deviation in $S_{,v}$ is a semiclassical effect), so the overall scaling goes as $M^{-3}$. We therefore check it by plotting this deviation multiplied by $M^{3},$and check that we get relatively constant line. This approximately constant line is seen in Figure VII.14 (note the narrow range of the vertical scale).
Figure VII.14: $\left(S_{,v}+\kappa_{-}\right)M^{3}$ on last u ray
@@
Subsequently, we show in Figure VII.15 the deviation from the analytical approximation for $R_{,v}$ for two different $\varepsilon_{0}$ values. This deviation is equal to $R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}(v)}{\kappa_{-}(v)}$, see Eq. (IV.40) .
@@
Figure VII.15: $R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ with two $\varepsilon_{0}$ values, one of them scaled by $\frac{1}{4}$, on the last u ray
@@
Figure VII.16: $\left(R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}\right)M^{3}$ scaling of the deviation on last u ray
We also wish to know how the mentioned deviation scales with $M$. This deviation scales as $\propto M^{-3}$, and we show it by plotting the deviation times $M^{3},$and check that we get a relatively constant curve. The reason that the deviation scales as $M^{-3}$ is because the dimensions of $R_{,v}]=[l]$ (as $R$ has dimensions of length squared), and we need to further multiply it by a factor of $\varepsilon_{0}^{2}\propto\frac{\hbar^{2}}{M^{4}}$, so that the total power of $M$ becomes $M^{-3}$. This is indeed seen in Figure VII.16 – note the small scale of the vertical axis.
@@
The following Figure VII.17 shows the deviation of $S_{,u}$ from $\kappa_{-}^{u}(u)$. Note that we’re now using $M^{u}(u)$ as the horizontal axis, as we need the mass function that depends on $u$. Like the case of $S_{,v}$, we expect the deviation of $\varepsilon_{0}=\frac{1}{500}$ to be twice as big as the deviation for $\varepsilon_{0}=\frac{1}{1000}$.
Figure VII.17: $S_{,u}+\kappa_{-}$ with two $\varepsilon_{0}$ values with one scaled by $\frac{1}{2}$, on the last v ray
We therefore scale the curve of $\varepsilon_{0}=\frac{1}{500}$ by a factor of $\frac{1}{2}$ in Figure VII.17 in green. This figure also shows the unscaled version in orange. The green and blue curves overlap, confirming the scaling by $\varepsilon_{0}$.
@@
VIII Numerical results – run with all sources set to 1
@@
VIII.1 Overview
Figure VIII.1: R on last u,v rays
@@ VIII.1 Overview
As you can observe in Figure VIII.1a, showing $R(v)$ along the last-$u$ ray, in region 1 $R$ starts at $r_{+}^{2}(M=1,Q=0.8)=2.56$, and slowly drifts down because of non-vanishing $M_{,v}$. Afterwards it quickly falls down during the active region (region 2) – just like in RN – and ends up near $r_{-}^{2}(M=1,Q=0.8)=0.16$ at the beginning of region 3. Up until region 3, this looks pretty much the same as the case of zero sources shown in Figure VII.1a.
However, in region 3, the two figures look differently: the difference in the value of the sources in these two cases leads to a difference in the corresponding values of $\widetilde{T}_{vv}^{(-)}$. In particular, the two cases have different signs for $\widetilde{T}_{vv}^{(-)}$. Thus, in Figure VIII.1a (run with non-vanishing sources) $R(v)$ goes up in region 3 because of negative $\widetilde{T}_{vv}^{(-)}$, whereas in Figure VII.1a (run with vanishing sources) $R$ goes down because of positive $\widetilde{T}_{vv}^{(-)}$. (Recall that $R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ and so $R_{,v}$ is of the opposite sign to $\widetilde{T}_{vv}^{(-)}$.)
In the next Figure VIII.1b, we look at $R(u)$ along the last $v$ ray. We can see that $R$ starts at $r_{+}^{2}=0.551535$ (corresponding to the final values of $M$ and $Q$ along the EH) and it is almost exactly constant throughout region 1, then it falls down in region 2, similar to Figure VII.1b. Because of the non-vanishing sources, $\widetilde{T}_{uu}$ no longer vanishes. In the present case, $\widetilde{T}_{uu}^{(-)}$ is negative and so $R(u)$ drifts upwards in region 3. This is in contrast to the case of Figure VII.1b where, due to vanishing sources, $\widetilde{T}_{uu}^{(-)}=0$ and $R(u)$ stays constant.
Figure VIII.2: $\widetilde{T}_{uu}(v)$ and $\widetilde{T}_{uu}(u)$ on last u,v rays
@@
The following Figure VIII.2 shows $\widetilde{T}_{uu}$ on the two rays, namely last-$u$ (Figure VIII.2a) and last-$v$ (Figure VIII.2b) rays. As can be seen in both panels, in region 1 near the initial hypersurface, $\widetilde{T}_{uu}=0$. Then the non-vanishing sources in region 2 cause a non-vanishing, negative $\widetilde{T}_{uu}$.
Figure VIII.3: $\widetilde{T}_{vv}(u,v)$
@@
In Figure VIII.3 we display $\widetilde{T}_{vv}$ on both rays, namely last-$u$ (Figure VIII.3a) and last-$v$ (Figure VIII.3b) rays. In Figure VIII.3a we can see that the chosen source values caused a negative $\widetilde{T}_{vv}$ in region 3, in contrast to Figure VII.4 where $\widetilde{T}_{vv}$ is positive there.
Plotting $\widetilde{T}_{vv}$ as a function of $u$ on the last-$v$ ray in Figure VIII.3b, we see that it is almost exactly constant in region 1 and 3, it only changes in region 2 as a result of the non-vanishing sources there.
Figure VIII.4: Scaling of $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$
@@
The scaling of both $\widetilde{T}_{vv}^{(+)}$ and $\widetilde{T}_{vv}^{(-)}$ should [^6] This simple scaling holds in our case because $\frac{Q}{M}$ is constant along the EH, and in addition all source terms were constructed with the appropriate $M$ scaling. be $\propto M^{-2}$ and to illustrate that, we plot $\widetilde{T}_{vv}M^{2}$ in Figure VIII.4a. This quantity is approximately constant in both region 1 and region 3, confirming that $\widetilde{T}_{vv}\propto M^{-2}$ there.
Next we do a similar check but for $\widetilde{T}_{uu}(u)$ rather than $\widetilde{T}_{vv}(v)$. To this end, we now plot $\widetilde{T}_{uu}M^{u}(u)^{2}$ as a function of $u$ in Figure VIII.4b. This quantity is again approximately constant in region 1 and region 3, confirming that $\widetilde{T}_{uu}\propto M^{u}(u)^{-2}$ in both these regions.
Because the sources only effect the metric during and after region 2, in region 1 the results are (almost) exactly the same as the no-sources run. In particular the behavior of $\widetilde{T}_{vv}(v)$ on the initial hypersurface in the present case is exactly the same as seen in Figure VII.6a.
VIII.2 Region 3: comparison with the analytical approximation
Like in the corresponding subsection VII.3, we will now compare $R_{,u},$$R_{,v},$$S_{,u}$$S_{,v}$ with their expected values according to Eqs. (IV.5,IV.16,IV.40,IV.41).
Figure VIII.5: Analytical approximation of $S_{,v}$ and $S_{,u}$

### v2

@@ 7.2 Region 2: comparison with the local RN approximation
Figure 7.7: computed R vs. RN value of R
@@
In Figure 7.7a we compare the numerically-computed $R$ with that corresponding to the RN metric (in double null Eddington coordinates) but with its $M$ and $Q$ parameters replaced with $M(v)$ and $Q^{v}(v)$. Notice the complete overlap of the two curves .
Subsequently, Figure 7.7b shows $R(u)$ along the last $v$ ray, versus RN with constant $M,Q$ corresponding to $M(v=900)$ and $Q^{v}(v=900)$. Notice that because there is no drift of $M$ and $Q$ parameters in the $u$ direction, the two curves nicely overlap.
Figure 7.8: S computed vs. RN
@@
We move on to show $S$ in Figure 7.8 in region 2, versus the value of $S$ in the corresponding RN metric, determined by
$\frac{e^{S}}{r}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (7.2)
In Panel 7.8a we plot $S(v)$ along the u=60u=-60 ray. Then in panel 7.8b we plot $S(u)$ along the last ray $v=900$. In both cases, the two curves are in good agreement. The visible deviation at late times is expected because this range is further from the domain of validity of the local RN approximation which is our focus here.
7.3 Region 3: comparison with the analytical approximation
In this subsection, we will compare the numerically computed quantities $R_{,u},$$R_{,v},$$S_{,u}$$S_{,v}$ in region 3 with their predicted values according to Eqs. (4.5,4.16,4.40,4.41).
Figure 7.9: Numerical result and analytical approximation for $S_{,v}$, $S_{,u}$
@@ 7.3 Region 3: comparison with the analytical approximation
We start with Figure 7.9a, which compares $S_{,v}$ and $\kappa_{-}(v)$, both displayed as a function of $v$. They are indeed very close, as predicted by the analytical approximation. Then we see in Figure 7.9b that the approximation is good for $S_{,u}$ as well, when comparing it to $\kappa_{-}^{u}$.
Figure 7.10: Analytical approximation of $R_{,v}$, $R_{,u}$ compared with their numerical values
@@
Moving on to the derivatives of $R$, we see in Figure 7.10a that the analytical approximation Eq. (4.40) works, comparing $R_{,v}$ with $\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ as they vary as functions of $v$.
Finally, we show in Figure 7.10b the quantities $R_{,u}$ and $\frac{\widetilde{T}_{uu}^{(-)}}{\kappa_{-}^{u}}$ in region 3. These quantities should both vanish, as mentioned in subsection 2.8, because we have no sources in the present run. In the figure we see that both quantities are very small. It is remarkable that their erroneous non-vanishing value is still consistent with Eq. (4.41). Our interpretation of this situation is as follows: these erroneous non-vanishing values happens because of finite-$du$ error. This causes a drift in $\widetilde{T}_{uu}$ leading to non-vanishing $\widetilde{T}_{uu}^{(-)}$– which in turn leads to non-vanishing $R_{,u}$ in region 3, in accordance with Eq. (4.41).
To check this interpretation, we show in Figure 7.11 the magnitude of $R_{,u}$ for two different values of $du$: $\frac{1}{100}$ and $\frac{1}{50}$. The third curve shows the $du=\frac{1}{50}$ case multiplied by $\frac{1}{4}$. We can see that it perfectly overlaps with the curve of $du=\frac{1}{100}$. This shows that the erroneous non-vanishing value of $R_{,u}$ indeed scales with $du^{2}$, implying that it is associated with a finite-$du$ truncation error.
Figure 7.11: $R_{,u}$ with different $du$ values on last v ray
@@
7.4 Deviation from the analytical approximation (region 3)
@@
First, we plot the deviation $S_{,v}+\kappa_{-}$ in Figure 7.12a. (To appreciate the magnitude of this deviation, note that it is roughly about 1% of the value of $S_{,v}$ as seen in Figure 7.9a.)
Figure 7.12: Deviation from analytical approximation of $S_{,v}$, $R_{,v}$
@@
We also show the deviation of $R_{,v}$ in Figure 7.12b
To analyze whether this deviation scales with $\varepsilon_{0},$we do two runs: one with $\varepsilon_{0}=\frac{1}{1000}$ as usual, and another one with $\varepsilon_{0}=\frac{1}{500}.$ To compare the two different runs, we plot the mentioned deviations as a function of $M^{v}(v)$ instead of as a function of $v$, to account for the fact that the evaporation rate is different in these two cases. Figure 7.13 compares two different $\varepsilon_{0}$ runs. (Note that in this plot, the low $v$ values correspond to high $M$, with $v=0$ at $M=1$.) If the deviation is proportional to $\varepsilon_{0}$, we would expect it to be twice as big in $\varepsilon_{0}=\frac{1}{500}$ compared to $\varepsilon_{0}=\frac{1}{1000}$.
Figure 7.13: $S_{,v}+\kappa_{-}$ with two $\varepsilon_{0}$ values, scaled by a factor of $\frac{1}{2}$, on the last u ray
To test this assumption, Figure 7.13 scales the curve of $\varepsilon_{0}=\frac{1}{500}$ by $\frac{1}{2}$ (green curve), and now we can see the two curves (green and blue) overlap, confirming that the deviation indeed scales with $\varepsilon_{0}$–meaning that it is a finite-$\varepsilon_{0}$ effect. We also display the unscaled $\varepsilon_{0}=\frac{1}{500}$ curve, as is, in orange. Note that the graph does not start from $M=1$, because here we focus on the analytical approximation of region 3, therefore we restricted the curves to the range of $v$ for which the last $u$ ray is within region 3. In the next figure we see that in earlier $M$ values there is a large deviation from the analytical approximation of region 3.
We also wish to know how this finite-$\varepsilon_{0}$ deviation scales with $M$. From basic semiclassical scaling, one would expect this deviation to scale as $\propto M^{-3}$. This is because the dimensions of $S_{,v}$ is $S_{,v}]=[l]^{-1}$ (as $S$ is dimensionless), and we need to multiply this by the dimensionless factor $\frac{\hbar}{M^{2}}$ (as this deviation in $S_{,v}$ is a semiclassical effect), so the overall scaling goes as $M^{-3}$. We therefore check it by plotting this deviation multiplied by $M^{3},$and check that we get relatively constant line. This approximately constant line is seen in Figure 7.14 (note the narrow range of the vertical scale).
Figure 7.14: $\left(S_{,v}+\kappa_{-}\right)M^{3}$ on last u ray
@@
Subsequently, we show in Figure 7.15 the deviation from the analytical approximation for $R_{,v}$ for two different $\varepsilon_{0}$ values. This deviation is equal to $R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}(v)}{\kappa_{-}(v)}$, see Eq. (4.40) .
@@
Figure 7.15: $R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ with two $\varepsilon_{0}$ values, one of them scaled by $\frac{1}{4}$, on the last u ray
@@
Figure 7.16: $\left(R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}\right)M^{3}$ scaling of the deviation on last u ray
We also wish to know how the mentioned deviation scales with $M$. This deviation scales as $\propto M^{-3}$, and we show it by plotting the deviation times $M^{3},$and check that we get a relatively constant curve. The reason that the deviation scales as $M^{-3}$ is because the dimensions of $R_{,v}]=[l]$ (as $R$ has dimensions of length squared), and we need to further multiply it by a factor of $\varepsilon_{0}^{2}\propto\frac{\hbar^{2}}{M^{4}}$, so that the total power of $M$ becomes $M^{-3}$. This is indeed seen in Figure 7.16 – note the small scale of the vertical axis.
@@
The following Figure 7.17 shows the deviation of $S_{,u}$ from $\kappa_{-}^{u}(u)$. Note that we’re now using $M^{u}(u)$ as the horizontal axis, as we need the mass function that depends on $u$. Like the case of $S_{,v}$, we expect the deviation of $\varepsilon_{0}=\frac{1}{500}$ to be twice as big as the deviation for $\varepsilon_{0}=\frac{1}{1000}$.
Figure 7.17: $S_{,u}+\kappa_{-}$ with two $\varepsilon_{0}$ values with one scaled by $\frac{1}{2}$, on the last v ray
We therefore scale the curve of $\varepsilon_{0}=\frac{1}{500}$ by a factor of $\frac{1}{2}$ in Figure 7.17 in green. This figure also shows the unscaled version in orange. The green and blue curves overlap, confirming the scaling by $\varepsilon_{0}$.
@@
8 Numerical results – run with all sources set to 1
@@
8.1 Overview
Figure 8.1: R on last u,v rays
@@ 8.1 Overview
As one can observe in Figure 8.1a, showing $R(v)$ along the last-$u$ ray, in region 1 $R$ starts at $r_{+}^{2}(M=1,Q=0.8)=2.56$, and slowly drifts down because of non-vanishing $M_{,v}$. Afterwards it quickly falls down during the active region (region 2) – just like in RN – and ends up near $r_{-}^{2}(M=1,Q=0.8)=0.16$ at the beginning of region 3. Up until region 3, this looks pretty much the same as the case of zero sources shown in Figure 7.1a.
However, in region 3, the two figures look differently: the difference in the value of the sources in these two cases leads to a difference in the corresponding values of $\widetilde{T}_{vv}^{(-)}$. In particular, the two cases have different signs for $\widetilde{T}_{vv}^{(-)}$. Thus, in Figure 8.1a (run with non-vanishing sources) $R(v)$ goes up in region 3 because of negative $\widetilde{T}_{vv}^{(-)}$, whereas in Figure 7.1a (run with vanishing sources) $R$ goes down because of positive $\widetilde{T}_{vv}^{(-)}$. (Recall that $R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ and so $R_{,v}$ is of the opposite sign to $\widetilde{T}_{vv}^{(-)}$.)
In the next Figure 8.1b, we look at $R(u)$ along the last $v$ ray. We can see that $R$ starts at $r_{+}^{2}=0.551535$ (corresponding to the final values of $M$ and $Q$ along the EH) and it is almost exactly constant throughout region 1, then it falls down in region 2, similar to Figure 7.1b. Because of the non-vanishing sources, $\widetilde{T}_{uu}$ no longer vanishes. In the present case, $\widetilde{T}_{uu}^{(-)}$ is negative and so $R(u)$ drifts upwards in region 3. This is in contrast to the case of Figure 7.1b where, due to vanishing sources, $\widetilde{T}_{uu}^{(-)}=0$ and $R(u)$ stays constant.
Figure 8.2: $\widetilde{T}_{uu}(v)$ and $\widetilde{T}_{uu}(u)$ on last u,v rays
@@
The following Figure 8.2 shows $\widetilde{T}_{uu}$ on the two rays, namely last-$u$ (Figure 8.2a) and last-$v$ (Figure 8.2b) rays. As can be seen in both panels, in region 1 near the initial hypersurface, $\widetilde{T}_{uu}=0$. Then the non-vanishing sources in region 2 cause a non-vanishing, negative $\widetilde{T}_{uu}$.
Figure 8.3: $\widetilde{T}_{vv}(u,v)$
@@
In Figure 8.3 we display $\widetilde{T}_{vv}$ on both rays, namely last-$u$ (Figure 8.3a) and last-$v$ (Figure 8.3b) rays. In Figure 8.3a we can see that the chosen source values caused a negative $\widetilde{T}_{vv}$ in region 3, in contrast to Figure 7.4 where $\widetilde{T}_{vv}$ is positive there.
Plotting $\widetilde{T}_{vv}$ as a function of $u$ on the last-$v$ ray in Figure 8.3b, we see that it is almost exactly constant in region 1 and 3, it only changes in region 2 as a result of the non-vanishing sources there.
Figure 8.4: Scaling of $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$
@@
The scaling of both $\widetilde{T}_{vv}^{(+)}$ and $\widetilde{T}_{vv}^{(-)}$ should [^7] This simple scaling holds in our case because $\frac{Q}{M}$ is constant along the EH, and in addition all source terms were constructed with the appropriate $M$ scaling. be $\propto M^{-2}$ and to illustrate that, we plot $\widetilde{T}_{vv}M^{2}$ in Figure 8.4a. This quantity is approximately constant in both region 1 and region 3, confirming that $\widetilde{T}_{vv}\propto M^{-2}$ there.
Next we do a similar check but for $\widetilde{T}_{uu}(u)$ rather than $\widetilde{T}_{vv}(v)$. To this end, we now plot $\widetilde{T}_{uu}M^{u}(u)^{2}$ as a function of $u$ in Figure 8.4b. This quantity is again approximately constant in region 1 and region 3, confirming that $\widetilde{T}_{uu}\propto M^{u}(u)^{-2}$ in both these regions.
Because the sources only affect the metric during and after region 2, in region 1 the results are (almost) exactly the same as the no-sources run. In particular the behavior of $\widetilde{T}_{vv}(v)$ on the initial hypersurface in the present case is exactly the same as seen in Figure 7.6a.
8.2 Region 3: comparison with the analytical approximation
Like in the corresponding subsection 7.3, we will now compare $R_{,u},$$R_{,v},$$S_{,u}$$S_{,v}$ with their expected values according to Eqs. (4.5,4.16,4.40,4.41).
Figure 8.5: Analytical approximation of $S_{,v}$ and $S_{,u}$